# Modeling Visual Memorability Assessment with Autoencoders Reveals Characteristics of Memorable Images

### v1

@@ Abstract
Background: Image memorability refers to the phenomenon where certain images are more likely to be remembered than others. It is a quantifiable and intrinsic image attribute, defined as the likelihood of being remembered upon a single exposure. Despite advances in understanding human visual perception and memory, it is unclear what features contribute to an image’s memorability. To address this question, we propose a deep learning-based computational modeling approach.
Methods: We modeled the subjective experience of visual memorability using an autoencoder based on VGG16 Convolutional Neural Networks (CNNs). The model was trained on images for one epoch, to simulate the single-exposure condition used in human memory tests. We investigated the relationship between memorability and reconstruction error, assessed latent space representations’ distinctiveness, and developed a Gated Recurrent Unit (GRU) model to predict memorability likelihood. Interpretability analysis was conducted to identify key image characteristics contributing to memorability.
Results: Our results demonstrate a significant correlation between the images’ memorability score and autoencoder’s reconstruction error, and the robust predictive performance of its latent representations. Distinctiveness in these representations correlated significantly with memorability. Additionally, certain visual characteristics, such as strong contrasts, distinctive objects, and prominent foreground elements were among the features contributing to image memorability in our model.
Conclusions: Images with unique features that challenge the autoencoder’s capacity are inherently more memorable. Moreover, these memorable images are distinct from others the model has encountered, and the latent space of the encoder contains features predictive of memorability.
@@ 1 Introduction
In this work, we develop a computational model to simulate subjective experiences in the visual memory experiment used to quantify image memorability. Unlike previous studies, our work emphasizes unsupervised learning, specifically autoencoders, aiming to investigate the potential of predicting memorability without relying on labeled datasets. We employ transfer learning with a pre-trained autoencoder based on Convolutional Neural Networks (CNNs), here VGG16 architecture, initially trained on the diverse ImageNet dataset. We fine-tune this model on a categorical memorability dataset, MemCat [16]. We then investigate the correlation between image reconstruction error, latent space distinctiveness, and memorability. We hypothesize that images with higher reconstruction errors—indicative of greater complexity or uniqueness—are inherently more memorable. Additionally, inspired by the visual memory game used to quantify image memorability, we explore the predictive power of the latent space representations by developing a Recurrent Neural Network model, GRU, to predict the likelihood of the images being remembered in the experiment. We hypothesize that encoded latent representations contain sufficient information to predict memorability scores. Finally, we utilize an explainable AI technique, Integrated Gradients (IG), to visually and quantitatively assess which specific image features contribute to memorability.
2 Methods
2.1 Proposed approach
@@ 2.2 Dataset
2.3 Model development
To simulate image memorability assessment, we used a VGG16-based autoencoder pre-trained on the ImageNet dataset [18, 19]. VGG16 is a well-established model in computer vision [20], commonly used in tasks such as image classification and style transfer [21]. It utilizes deep convolutional neural networks (CNNs), which are loosely inspired by how the human visual cortex processes images. CNNs like VGG16 perform hierarchical feature extraction, starting with simple features such as edges and gradually capturing more complex patterns. The deep convolutional layers of VGG16 allow it to extract rich and efficient features from images, which aligns with our objective of understanding image memorability. The pre-trained VGG16 model simulates the accumulated visual experience of human participants.
@@ 2.3 Model development
2.3.1 Modeling memorability experiment
The autoencoder was fine-tuned on the MemCat dataset for a single epoch with a batch size of one, mimicking the human memorability experiment paradigm, where memorability is quantified by the likelihood of remembering an image after participants view images only once. The single-exposure condition, commonly employed in human memorability assessments, was thus applied to evaluate if a pre-trained autoencoder could be used to mimic human-like memorability assessment based on a single exposure to images.
@@ 2.3.1 Modeling memorability experiment
Figure 2 demonstrates the schematic representation of our approach for modeling memorability assessment experiment by an autoencoder.
2.4 Memorability prediction using autoencoder latent space representations
To explore whether the latent code of the autoencoder could be used to predict memorability, we utilized the encoder part of our fine-tuned autoencoder to extract latent codes from the MemCat dataset. Recognizing the potential sequential nature of how image features impact memorability, we employed Recurrent Neural Networks (RNNs) for this task, specifically training a Gated Recurrent Unit (GRU) model [28]. GRUs are an efficient type of RNN designed to capture temporal dependencies in data, making them suitable for this analysis. The GRU model was trained on the sequence of latent codes generated by the encoder. The dataset was split into training, validation, and test sets, with 5892 images, 931 images, and 3177 images, respectively, all maintaining a mean memorability score of approximately 0.693. Hyperparameter optimization was conducted, testing learning rates from 0.0001 to 0.1, GRU sequence lengths from 2 to 20, hidden layer sizes of 256, 512, and 1024, varying the number of hidden layers from 1 to 3, and using a dropout ratio from 0 to 0.5. Early stopping with patience of 10 was employed to determine the optimal number of training epochs.
The goal of this analysis was to determine whether memorability behavior could be predicted solely from the latent representations of images in a sequential paradigm similar to the memorability paradigm used with human participants. A successful prediction would imply that these latent representations capture essential characteristics influencing memorability, validating the autoencoder’s effectiveness in generating cognitively relevant feature representations.
2.5 Model interpretation
@@
To better understand what image features contribute to memorability prediction using the latent space representations, we employed Integrated Gradients (IG), an Explainable AI (XAI) technique. IG attributes the prediction of a neural network to its input features by computing the gradients of the output with respect to the input at points interpolated between a given baseline and the actual input [29]. This method is particularly suited for our complex architecture, which includes an autoencoder followed by a GRU model.
@@
For IG analysis, the baseline input was set to be an image created by averaging all training set images, representing the absence of distinctive features. IG attributes were computed by averaging the gradients over 50 steps between the baseline and the original image. These gradients provided a detailed map of feature contributions, helping us identify which aspects of the image were most influential in making it memorable according to the model’s assessment. The analysis was conducted using the Captum library [30] in PyTorch.
Figure 2: Schematic representation of the approach to model memorability assessment experiment. The autoencoder, pre-trained on ImageNet, is fine-tuned on the MemCat image memorability dataset for one epoch and batch size of one.
3 Results
3.1 Memorable images are harder to reconstruct and more distinct
@@
To evaluate whether there is a difference in the model’s performance in reconstructing images with varying memorability, we calculated Spearman’s rank correlation between the images’ memorability scores and their reconstruction errors. This reconstruction error served as a quantitative metric to assess the model’s accuracy in reproducing the input data. To ensure comparability across different loss functions, the reconstruction errors were normalized to fall within the range of [0,1].
@@
Our results show a significant positive correlation between the autoencoder’s reconstruction error and memorability scores. The model fine-tuned using the LPIPS (SqueezeNet) loss function with a learning rate of $065E^{-}$ 06 yielded the highest correlation between reconstruction error and memorability scores. This model, fine-tuned on all MemCat images, was selected for the remainder of the analysis in this paper. The Spearman’s rank correlation for this selected model was 0.452, (CI: 0.442, 0.467), with a p-value approaching zero.
@@
Furthermore, we investigated the distinctiveness of images within the autoencoder’s latent space, hypothesizing that more distinctive images are more likely to be memorable. Distinctiveness in memorability research is typically defined as the Euclidean distance between a target image and its nearest neighbor [15]. However, this measure can be sensitive to the dataset’s size. To address this, we introduced a normalized metric, standardizing the distances to the nearest neighbor using Z-score normalization. This provides a robust, dataset-independent metric of distinctiveness and ensures that distinctiveness remains robust across different dataset sizes, providing a stable and consistent measure of how unique each image is within its context.
@@
This different behavior in food images might be due to having higher memorability scores than the rest of the images in the MemCat dataset. We observe in Figure 1 that the memorability scores of food images are concentrated in the very high memorable range which may impact the model’s performance on them.
@@
3.2 The memorability effect diminishes with multiple exposures
To examine how continued fine-tuning affects the observed correlation between model reconstruction error and memorability, the autoencoder was trained for up to 20 epochs. As expected, the correlation between reconstruction error and memorability decreased as the model was exposed to the images multiple times. After 20 epochs, the correlation dropped significantly to 0.22, (CI: 0.21, 0.24), with $p-value=2.8\times 10^{-}114$, suggesting that repeated exposures to an image reduce its impact on memorability.
3.3 Models trained with perceptual loss show a stronger effect
@@
3.4 Memorability behavior can be predicted from the autoencoder latent representation
@@
According to the results, our memorability prediction model demonstrated robust performance in predicting memorability from the latent representations generated by the encoder. On the test set, the GRU model achieved an area under the ROC curve (AUC) of 0.72 (CI: 0.694, 0.744) and a Precision-Recall (PR) AUC of 0.65 (CI: 0.612, 0.682). Additionally, the model’s predicted scores correlated significantly with the original memorability scores with correlation = 0.47, (CI: 0.431, 0.505), $p-value=1.5\times 10^{-87}$.
The best-performing GRU model consisted of two hidden layers, each with 512 neurons, and used a sequence length of 2. The learning rate was 0.0001, with a batch size of 16. The model was trained for 26 epochs with early stopping with patience of 10. The total training and validation time per run was approximately 20 minutes.
@@
Furthermore, the classification of images into high and low memorability groups based on a median score of 0.74 yielded an accuracy of 0.67 (CI: 0.66, 0.71), precision of 0.64 (CI: 0.60, 0.68), sensitivity of 0.64 (CI: 0.66, 0.72), specificity of 0.71 (CI: 0.63, 0.69), and an F1 score of 0.65 (CI: 0.62, 0.68).
3.5 Model Interpretation identifies image characteristics contributing to memorability
@@
We employed the Integrated Gradients (IG) method to interpret the prediction model’s predictions and visually assess which regions of an image were most influential in determining memorability. Figure 5 provides visual explanations of IG attributes for selected images from the MemCat dataset, highlighting regions with the highest attribution values.
Distinctive features such as bright colors, unique shapes, and recognizable objects (e.g., the lion’s face, yellow bell pepper) were crucial for memorability. Faces, human figures, and interaction scenes also significantly contributed to memorability.
Figure 5: Integrated Gradients attributes for sample images from the MemCat dataset. The numbers on top of the images show the Memorability score, in descending order from the top left.
4 Discussion

### v2

@@ Abstract
Image memorability refers to the phenomenon where certain images are more likely to be remembered than others. It is a quantifiable and intrinsic image attribute, defined as the likelihood of an image being remembered upon a single exposure. Despite advances in understanding human visual perception and memory, it is unclear what features contribute to an image’s memorability. To address this question, we propose a deep learning-based computational modeling approach. We employ an autoencoder-based approach built on VGG16 convolutional neural networks (CNNs) to learn latent representations of images. The model is trained in a single-epoch setting, mirroring human memory experiments that assess recall after a single exposure. We examine the relationship between autoencoder reconstruction error and memorability, analyze the distinctiveness of latent space representations, and develop a multi-layer perceptron (MLP) model for memorability prediction. Additionally, we perform interpretability analysis using Integrated Gradients (IG) to identify the key visual characteristics that contribute to memorability. Our results demonstrate a significant correlation between the images’ memorability score and the autoencoder’s reconstruction error, as well as the robust predictive performance of its latent representations. Distinctiveness in these representations correlated significantly with memorability. Additionally, certain visual characteristics were identified as features contributing to image memorability in our model. These findings suggest that autoencoder-based representations capture fundamental aspects of image memorability, providing new insights into the computational modeling of human visual memory.
@@ Introduction
In this work, we develop a computational framework to simulate subjective experiences in the visual memory experiment used to quantify image memorability. Unlike previous studies that focus on supervised memorability prediction, our approach emphasizes unsupervised learning through autoencoders, investigating whether image memorability can be inferred from reconstruction-based representations without relying on labeled datasets. We employ transfer learning using a VGG16-based autoencoder, pre-trained on the large-scale ImageNet dataset and fine-tuned on the categorical MemCat dataset [16]. To understand the relationship between image structure and memorability, we analyze image reconstruction error and latent space distinctiveness, hypothesizing that higher reconstruction errors—indicative of greater complexity or uniqueness—correlate with higher memorability. We further assess the predictive capacity of latent space representations by training a multi-layer perceptron to estimate memorability scores. We hypothesize that these learned representations encode necessary visual information to distinguish between memorable and less memorable images. To identify the specific features contributing to memorability, we employ Integrated Gradients (IG), an explainable AI technique that quantitatively and visually highlights the most influential image attributes.
While many models have been developed to predict memorability scores from images, the goal of this study is not to propose a new predictive framework. Instead, we leverage an autoencoder to analyze what makes images memorable by examining reconstruction errors, latent space distinctiveness, and feature representations. Our findings contribute to both computational modeling of memory processes and practical applications in content creation, advertising, and education, where designing memorable visuals is of critical importance.
Methods
Proposed approach
@@ Methods
Model development
To simulate image memorability assessment, we used a CNN-based (VGG16) autoencoder pre-trained on the ImageNet dataset [18, 19]. VGG16 is a well-established model in computer vision [20], commonly used in tasks such as image classification and style transfer [21]. It utilizes deep convolutional neural networks (CNNs), which are loosely inspired by how the human visual cortex processes images. CNNs like VGG16 perform hierarchical feature extraction, starting with simple features such as edges and gradually capturing more complex patterns. The deep convolutional layers of VGG16 allow it to extract rich and efficient features from images, which aligns with our objective of understanding image memorability. The pre-trained VGG16 model simulates the accumulated visual experience of human participants.
@@
Modeling memorability experiment
The autoencoder was fine-tuned on the MemCat dataset for a single epoch, mimicking the human memorability experiment paradigm, where memorability is quantified by the likelihood of remembering an image after participants view images only once. The single-exposure condition, commonly employed in human memorability assessments, was thus applied to evaluate if a pre-trained autoencoder could be used to mimic human-like memorability assessment based on a single exposure to images.
@@
Memorability prediction using autoencoder latent space representations
To investigate the potential of autoencoder latent representations for predicting image memorability, we utilized the encoder component of our fine-tuned autoencoder to extract feature embeddings from the MemCat dataset. These embeddings represent a condensed, high-level abstraction of the input images, which we hypothesized to encapsulate cognitive and perceptual factors influencing memorability.
We employed a MLP with one to three hidden layers as predictive models for memorability classification. The models were designed to refine feature mappings from the pre-trained encoder output, capturing nonlinear relationships in the latent space. 1 to 3 hidden layers of 256 to 2048 neurons were tested.
A hyperparameter search was conducted to optimize model performance, evaluating learning rates between 0.01 and 0.00001 and batch sizes of 8, 16, 32, and 64. The models were trained using the AdamW optimizer, which incorporates decoupled weight decay for better generalization. To prevent overfitting, weight decay (L2 regularization) values of 1e-5, 1e-4, and 0.01 were explored. Additionally, dropout (ranging from 0.1 to 0.5) and layer normalization were incorporated to stabilize training and mitigate covariate shift. To ensure training efficiency and prevent overfitting, early stopping was applied with a patience of five epochs, halting training when validation loss stopped improving. These architectural and optimization choices ensured that the model effectively leveraged the encoder’s latent representations while maintaining robustness and avoiding overfitting.
The dataset was partitioned into training, validation, and test subsets, consisting of 5892, 931, and 3177 images, respectively, ensuring a balanced distribution with a mean memorability score of approximately 0.693.
The primary objective of this approach was to assess the capacity of the autoencoder’s latent space to encode cognitively relevant features predictive of memorability. Success in this endeavor would indicate that these latent codes effectively capture high-level visual and semantic information critical to memorability, thereby validating the encoder’s ability to generate task-relevant representations and shedding light on the inherent characteristics contributing to image memorability.
Explainability Analysis
@@
To better understand what image features contribute to memorability prediction using the latent space representations, we employed Integrated Gradients (IG), an Explainable AI (XAI) technique that attributes the prediction of a neural network to its input features by computing the gradients of the output with respect to the input at points interpolated between a given baseline and the actual input [28]. This method is particularly suited for our complex architecture, which includes an autoencoder followed by a feed-forward model.
@@
The baseline input was set as a blurred mean image, computed by first averaging all training images and then applying a Gaussian blur (kernel size = 15). This ensures a smooth baseline representation that retains overall structure but removes fine details. IG attributes were computed by averaging the gradients over 50 steps between the baseline and the original image. These gradients provided a detailed map of feature contributions, helping us identify which aspects of the image were most influential in making it memorable according to the model’s assessment. The analysis was conducted using the Captum library [29] in PyTorch.
Feature Analysis
We further explored which high-level image features contributed to the model’s prediction of higher memorability. The features we considered included semantic, structural, color-based, and texture-related properties. These features were analyzed by correlating their values with both memorability scores and Integrated Gradients (IG) attributions, allowing us to assess their influence on the model’s predictions.
A pre-trained ResNet-50 model was used to extract deep features, summarizing the high-level visual information present in the images [30]. Specific metrics calculated included the number of distinct classes detected in each image, the total number of objects, and the average object size, computed as the mean area covered by detected objects. These measures provided insights into scene complexity and object diversity, which influence memorability by affecting perceptual load and distinctiveness.
For segmentation and object detection, we employed pre-trained DeepLabV3 and Faster R-CNN models [31, 32] to quantify object characteristics and scene complexity. The Separation Score, derived from DeepLabV3, measured how well objects stood out from the background, capturing foreground-background contrast, which can influence visual focus and memorability. Meanwhile, the Faster R-CNN model detected objects in the image, providing metrics such as the number of objects and their average size, which further informed the role of object richness in memorability.
Color features were extracted using the CIE-LAB and HSV color spaces, which better align with human perception. We computed mean values of the L (Lightness), A (green-red), and B (blue-yellow) channels to assess dominant color tones [33]. Saturation, representing color intensity, was measured in the HSV color space, as more vivid images tend to capture attention and enhance retention. Color diversity, computed using Shannon entropy, quantified the variation in color composition, reflecting the presence of multiple hues and measuring visual richness.
To analyze texture and structural complexity, we computed image entropy, which quantifies the randomness in pixel intensities, reflecting the information content in an image [34]. Texture energy, derived from the Gray Level Co-occurrence Matrix (GLCM), captured local variations in surface details, providing insight into the richness of textures [35]. The clutter score, calculated using edge detection via the Canny algorithm [36], quantified the density of structural details, as cluttered images require greater cognitive processing.
Each feature’s impact on image memorability was analyzed by correlating its values with average IG attributions from our trained model. This allowed us to determine which visual properties were most influential in shaping the model’s predictions and how these features aligned with human memory recall patterns.
Results
Memorable images are harder to reconstruct and more distinct
@@ Results
To evaluate whether there is a difference in the model’s performance in reconstructing images with varying memorability, we calculated Spearman’s rank correlation between the images’ memorability scores and their reconstruction errors. This reconstruction error served as a quantitative metric to assess the model’s accuracy in reproducing the input data. To ensure comparability across different loss functions, the reconstruction errors were normalized within each model to fall within the range of [0,1].
@@
Our results show a significant positive correlation between the autoencoder’s reconstruction error and memorability scores. The model fine-tuned using the LPIPS (SqueezeNet) loss function with a learning rate of $065E^{-}$ 06 achieved the highest correlation. This model, fine-tuned on all MemCat images, was selected for the remainder of the analysis in this paper. The Spearman’s rank correlation for this selected model was 0.452, (CI: [0.442, 0.467]), with a p-value approaching zero.
@@
Furthermore, we investigated the distinctiveness of images within the autoencoder’s latent space, hypothesizing that more distinctive images are more likely to be memorable. Distinctiveness in memorability research is typically defined as the Euclidean distance between a target image and its nearest neighbor [15]. However, this measure can be sensitive to the dataset’s size. To address this, we introduced a normalized metric, standardizing the distances to the nearest neighbor using Z-score standardization. This provides a robust, dataset-independent metric of distinctiveness and ensures that distinctiveness remains robust across different dataset sizes, providing a stable and consistent measure of how unique each image is within its context.
@@
In the reconstruction error analysis (Figure 2), the variability of reconstruction errors across memorability bins for food images is lower compared to other categories. While other categories show a clear increase in reconstruction error with memorability, the food category exhibits a more uniform distribution, indicating that the autoencoder’s reconstruction capability is less sensitive to variations in food image memorability. This could be attributed to the visual homogeneity of food images in terms of color, texture, and composition, which may limit the model’s ability to distinguish between memorable and less memorable images.
Similarly, in the distinctiveness analysis (Figure 3), the food category shows a narrower range of distinctiveness scores across memorability bins. Although distinctiveness generally increases with memorability, this trend is subdued for food images. The limited variability in distinctiveness scores may stem from the memorability scores of food images being concentrated within a narrow range, as observed in Figure 1. This concentration may reduce the relative distinctiveness of food images within the dataset.
Overall, the food category highlights the importance of category-specific features in determining memorability. The uniformity in visual characteristics of food images appears to weaken the correlation between memorability and both reconstruction error and distinctiveness. These findings suggest that the memorability of food images may rely on subtler cues that are not effectively captured by the current modeling approach. Future studies could explore alternative feature representations or loss functions tailored to homogeneous categories like food to enhance the model’s sensitivity to variations in memorability.
@@
The memorability effect diminishes with multiple exposures
To examine how continued fine-tuning affects the observed correlation between model reconstruction error and memorability, the autoencoder was trained for up to 20 epochs. As expected, the correlation between reconstruction error and memorability decreased as the model was exposed to the images multiple times. After 20 epochs, the correlation dropped significantly to 0.22, (CI: 0.21, 0.24), with $p-value=2.8\times 10^{-}114$, suggesting that repeated exposures to an image reduce its impact on memorability. This trend aligns with human memory studies, where repeated exposure leads to increased familiarity but reduced novelty, ultimately affecting recall performance. The model’s decreasing correlation with memorability across epochs suggests that its learned representations gradually shift towards more generalizable features rather than the initial distinctive patterns that drive single-exposure memorability.
Models trained with perceptual loss show a stronger effect
@@
Memorability behavior can be predicted from the autoencoder latent representation
The best-performing model was selected based on performance on the validation set. The best prediction model was a deep neural network with two hidden layers of 1024 and 512 neurons, trained with a learning rate of $5\times 10^{-5}$, a batch size of 8, a dropout rate of 0.3, and layer normalization. The model was optimized using AdamW with a weight decay of 0.01 and trained with early stopping using a patience of 5 epochs. The model was trained for 23 epochs.
@@
According to the results, our memorability prediction model demonstrated robust performance in predicting memorability from the latent representations generated by the encoder. For binary classification of memorability, on the unseen test set,the model achieved an area under the ROC curve (AUC) of 0.76 (CI: [0.74, 0.78]) and a Precision-Recall (PR) AUC of 0.74 (CI: [0.71, 0.76]). Additionally, the model’s predicted scores correlated significantly with the original memorability scores with correlation = 0.56 (CI: [0.53, 0.58], $p=2.21\times 10^{-258}$), supporting the effectiveness of latent representations in capturing memorability-related features.
@@
The binary classification of images into high and low memorability, with the threshold of the median memorability score, the model achieved an accuracy of 0.69 (CI: [0.68, 0.71]), a precision of 0.69 (CI: [0.67, 0.71]), a sensitivity of 0.70 (CI: [0.67, 0.72]), a specificity of 0.69 (CI: [0.67, 0.71]), and an F1 score of 0.69 (CI: [0.67, 0.71]).
Model Interpretation identifies image characteristics contributing to memorability
@@
We employed the Integrated Gradients (IG) method to interpret the predictions of our memorability model, visualizing which regions in an image contributed most to its memorability score. Figure 4 illustrates the IG attributions for selected images from the MemCat dataset, where red regions indicate areas with higher attribution values. These regions are identified as the most influential for the model’s prediction of memorability.
In the animal category, such as images of the rabbit, lion, and rooster, the IG heatmaps emphasize facial regions, particularly the eyes and prominent features like fur patterns. These observations suggest that distinct facial features and textures play a crucial role in making animals memorable. For example, the lion’s mane and the rooster’s sharp contours further underscore the importance of high-contrast and well-defined features in driving memorability.
For food-related images, including yogurt, wine, coffee, and hot dogs, the IG attributions concentrate on vivid textures, strong contrasts, and central elements. The smooth surface of the wine glass and the detailed textures of the food toppings are highly attributed. This highlights the role of vibrant colors, intricate patterns, and well-composed central features in influencing memorability in food images.
In landscape scenes, such as mountains, geysers, and vineyards, the IG heatmaps show strong attributions for horizon lines, separation between foreground and background, and specific details like trees or water surfaces. The prominence of natural structures and high-contrast boundaries suggests that such elements contribute significantly to the memorability of landscape images.
For sports and activity images, such as basketball players, rowers, and runners, the IG attributions highlight human figures and interactions. The focus on human faces, gestures, and motion-related elements, such as the oars and leg movements, underscores the model’s sensitivity to dynamic, socially relevant features. The emphasis on human-related elements demonstrates their intrinsic link to memorability.
In the vehicle category, including airplanes, trucks, and machinery, the IG maps emphasize distinct shapes and structural details, such as wheels, propellers, and edges of the vehicle. These findings suggest that clear geometric patterns and mechanical features enhance memorability in vehicle-related imagery.
Overall, the IG analysis reveals that the model consistently focuses on distinct, high-contrast regions, central objects, and unique features across all categories. Elements like bright colors, textures, and recognizable structures (e.g., faces, objects) emerge as critical contributors to memorability. Furthermore, the attribution maps highlight the role of image composition, with the arrangement of features and their saliency playing an essential role in drawing attention and enhancing recall. Notably, the distribution of attributions differs slightly between categories. While human-related and animal features often dominate their respective categories, natural landscapes and vehicles rely more on structural complexity and contrast. This variability suggests that memorability is influenced not only by individual features but also by the contextual arrangement and saliency of elements within an image.
Figure 4: Integrated gradient attributes for sample images from the MemCat dataset, showing regions influencing the memorability predictor outputs. The numbers on top of the images show the Memorability score. Red areas indicate strong contributions to the model’s predictions. Each row corresponds to one of the five categories: animals, food, landscapes, sports, and vehicles, with images arranged in ascending order of memorability scores.
Feature Importance in Predicting Memorability from Latent Representations
The feature importance analysis in this study provides a detailed understanding of which image attributes significantly contribute to memorability. Using correlations with Integrated Gradient (IG) attributions and original memorability scores, we identified several features from the MemCat dataset that aligned well with human-perceived memorability. The features having significant correlation values with original memorability scores and IG attributes are listed in Table 2. The p-values indicate the statistical significance of these correlations. These findings provide insights into the visual attributes that influence image memorability and validate the relevance of the autoencoder-based latent representations used in the prediction model.
The results indicate that object presence and arrangement strongly affect memorability. Images containing a greater number of distinct classes, a higher number of objects, and larger objects tend to be more memorable. This suggests that increased scene complexity and diversity provide more reference points for memory encoding. Larger objects likely draw more visual attention, while distinct elements within an image may facilitate richer encoding and longer retention.
Visual complexity also plays a critical role. The clutter score, derived from edge detection, quantifies the presence of structural details in an image. The positive correlation with memorability suggests that denser, detail-rich images require greater cognitive processing, making them more likely to be retained. Image entropy, which quantifies randomness in pixel intensities, further supports this observation. Higher entropy values indicate greater visual information content, reinforcing the idea that diverse textures and non-uniform compositions enhance memorability.
Color properties contribute significantly to image retention. The strong correlation between memorability and the LAB color channels suggests that color balance across green-red (A) and blue-yellow (B) axes influences memorability. Higher values along these axes indicate distinct and contrasting color compositions, which are more likely to capture attention. Higher saturation was also positively correlated with memorability, confirming that vivid and intense colors enhance image recall. Conversely, the negative correlation with mean color LAB L (lightness) suggests that excessively bright images may reduce contrast, making objects within them less distinguishable and, consequently, less memorable.
The separation score, measuring foreground-background distinction, was positively associated with memorability. This suggests that images where the main subject stands out clearly from the background are more likely to be remembered, likely due to enhanced object recognition and reduced perceptual competition.
Texture-related features exhibited mixed effects. Texture energy, which reflects uniformity in patterns, was negatively correlated with memorability, indicating that highly uniform textures may lack distinctiveness and fail to engage visual processing. Higher levels of structural complexity, captured through clutter and edge density, were positively correlated, suggesting that diverse and complex textures create additional memory cues.
The comparison between IG attributions and original memorability scores revealed a consistent trend in feature importance. Features that strongly correlated with memorability scores, such as the number of distinct objects, object size, and separation score, also exhibited high correlations with IG attributions. This suggests that the model’s predictions align with human perception of memorability, validating the autoencoder-based latent representations as effective for memorability modeling.
While most features maintained similar correlations across both measures, some variations were observed. Saturation and image entropy exhibited stronger correlations with IG attributions than with original memorability scores, indicating that the model assigns more importance to these properties during prediction than humans may consciously recognize. Conversely, some structural complexity measures, such as texture energy, had weaker correlations with IG attributions, suggesting that although these features contribute to memorability, they are not directly emphasized by the model in prediction tasks.
Overall, these findings highlight that a combination of object properties, color composition, structural complexity, and foreground-background separation influences memorability. The alignment between IG attributions and original memorability scores suggests that the model effectively captures the same visual attributes that determine human memory retention.
Table 2: Spearman’s rank correlation of image features with integrated gradients attributes and original memorability scores, including confidence intervals and p-values.
Feature Integrated Gradients Attributes Memorability Scores
Corr. 95% CI p-val. Corr. 95% CI p-val.
Mean color LAB A 0.585 [0.557, 0.608] $6.47\times 10^{-291}$ 0.347 [0.314, 0.377] $1.27\times 10^{-90}$
Distinct object categories 0.531 [0.506, 0.556] $1.15\times 10^{-230}$ 0.440 [0.411, 0.471] $1.32\times 10^{-150}$
Num. objects 0.463 [0.437, 0.489] $9.79\times 10^{-169}$ 0.350 [0.316, 0.380] $5.88\times 10^{-92}$
Mean color LAB B 0.418 [0.390, 0.450] $2.10\times 10^{-134}$ 0.228 [0.194, 0.260] $9.32\times 10^{-39}$
Saturation 0.413 [0.384, 0.445] $2.05\times 10^{-131}$ 0.157 [0.122, 0.191] $6.59\times 10^{-19}$
Average object size 0.404 [0.374, 0.432] $9.29\times 10^{-125}$ 0.480 [0.449, 0.510] $3.63\times 10^{-183}$
Separation score 0.321 [0.287, 0.351] $6.69\times 10^{-77}$ 0.273 [0.241, 0.305] $1.91\times 10^{-55}$
Color diversity 0.276 [0.244, 0.308] $1.04\times 10^{-56}$ 0.115 [0.084, 0.154] $9.42\times 10^{-11}$
Image entropy 0.160 [0.128, 0.193] $1.21\times 10^{-19}$ 0.102 [0.066, 0.134] $7.36\times 10^{-9}$
Clutter score 0.136 [0.102, 0.169] $1.35\times 10^{-14}$ 0.124 [0.086, 0.158] $2.79\times 10^{-12}$
Mean color LAB L -0.123 [-0.156, -0.088] $3.61\times 10^{-12}$0.058 [-0.092, -0.022] $9.82\times 10^{-4}$
Texture energy -0.108 [-0.140, -0.075] $1.10\times 10^{-9}$0.078 [-0.113, -0.046] $9.57\times 10^{-6}$
Discussion
We observed a strong positive correlation between the model reconstruction error of images and image memorability scores, supporting the hypothesis that more memorable images are inherently more challenging for the autoencoder to reconstruct. Prior works have modeled memorability using classification models or generative approaches [37, 15]. In contrast, our study focuses on leveraging autoencoder-based feature representations to understand why certain images are harder to reconstruct and how this relates to memorability. Further, our results showed that the distinctiveness of images, as captured by the autoencoder’s latent space, also correlates with memorability. This strong positive correlation indicates that images with more pronounced features, represented by higher normalized distances from their nearest neighbors, are more memorable. This suggests that memorability may not be solely dependent on the presence of specific objects or textures but emerges from the structural and compositional uniqueness of an image, making it more difficult for the model to encode efficiently.